OpenStudy (anonymous):

log3(x+2)+log3(x+1)=2 exact solution...

OpenStudy (anonymous):

you know the properties of logs? you can write the expression on the left side as 1 logarithm... can you do that?

OpenStudy (anonymous):

I have it as log(3x+6)+log(3(x+1))-2

OpenStudy (anonymous):

use this property: \(\large log_bM + log_bN=log_b(M\cdot N) \)

OpenStudy (anonymous):

wait.. what is this the problem: \(\large log3(x+2)+log3(x+1)=2 \) or \(\large log_3(x+2)+log_3(x+1)=2 \)

OpenStudy (anonymous):

log3(x+2)+log3(x+1)=2

OpenStudy (anonymous):

so this is log base 10 ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok... the property still applies no matter what base we're talking.... can you write out the expression on the left as ONE logarithm (with base 10)???

OpenStudy (anonymous):

I don't know how to do this problem, the answer I am getting is x=-3/2+1/6 =sqrt 409 I need exact solution and move decimel 3

OpenStudy (anonymous):

how are you getting that answer (if you don't know how to do this problem?)

OpenStudy (anonymous):

I have it set up as log(3x+6)+log(3(x+1))-2 am I not setting it up write? My answer don't seem right

OpenStudy (anonymous):

since you said this is log base 10, then \[\large log3(x+2)+log3(x+1)=log[3\cdot 3\cdot (x+2)\cdot (x+1)] \] \[\large =log[9\cdot (x^2+3x+2)] \] \[\large =log(9x^2+27x+18) \] so now your equation looks like this: \(\large log(9x^2+27x+18)=2 \)

OpenStudy (anonymous):

Ok, my answer is 100?